OpenStudy (didierr):

If the product of 14 and a number is decreased by 7, and then the result is divided by 5, the answer is 7. What is the unknown number?

OpenStudy (photon336):

the product of a 14 and a number. Product means multiply so 14n is decreased by 7 we subtract 7 14n-7 divided by 5 \[\frac{ 14n-7 }{ 5 } = 7 \] multiply both sides by 5 \[\frac{ 14n-7 }{ 5 }*5 = 7 *5 \] then add 7 to both sides \[\frac{ 14n-7+7 }{ } = 35+7\] \[14n = 42 \] divide both sides by 14 \[\frac{ 42 }{ 14 } = 3 = n \] n = 3

OpenStudy (photon336):

@didierr